3x^2+2y^2=30
2x^2-5y^2=-37

Answer by Alan3354(69443)

3x^2+2y^2=30
2x^2-5y^2=-37
-----------------
If you want to solve for x & y, you should say that.
---
6x^2 + 4y^2 = 60 --- 1st eqn times 2
6x^2 -15y^2 = -111 --- 2nd eqn times 3
--------------------------------------------- Subtract
19y^2 = 171
y^2 = 9
y = +3, -3
==================
Sub for y to find values of x.